what is the equatio for a line perpendicular to 3x+4y=12 and passes through (7,1)?
thank you :)

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
perpendicular to 3x+4y=12 and passes through (7,1)?
3x+4y=12 y= (-3/4)x + 3 ( m = (-3/4))
NEw Line: Perpendicular Line will have slope that is the negative reciprocal,
y = (4/3)x + b |Using ordered pair (7,1) to solve for b
1 = (4/3)*7 + b
-25/3 = b
y = (4/3)*x - 25/3